Kiting during a storm. The legend that Benjamin Franklin flew a kite as a storm approached is only a legend — he was neither stupid nor suicidal. Suppose a kite string of radius 2.11 mm extends directly upward by 0.815 km and is coated with a 0.520 mm layer of water having resistivity 178 ?·m. If the potential difference between the two ends of the string is 168 MV, what is the current through the water layer? The danger is not this current but the chance that the string draws a lightning strike, which can have a current as large as 500 000 A (way beyond just being lethal).Chapter 26, Problem 022 Flying Circus of Physics Kiting during a storm. The legend that Benjamin Franklin flew a kite as a storm approached is only a legend-he was neither stupid nor suicidal. Suppose a kite string of radius 2.11 mm extends directly upward by 0.815 km and is coated with a 0.520 mm ayer of water having resistivity 178 Ω m. If the potential difference between the two ends of the string 168 what st e current through the water layer? The danger is not this current but the chance that the string draws a lightning strike, which can have a current as large as 500 000 A (way beyond just being lethal) Number Units
Answer #1

area of cross section

A = 3.14*(((2.11+0.52)*10^-3)^2-(2.11*10^-3)^2) = 7.74*10^-6 m^2

resistance of the wire R = rho*l/A

R = 178*0.815*10^3/7.74*10^-6

R = 18.74*10^9 ohm

current through the wire is

i = v/R

i = 168*10^6/18.74*10^9

i = 8.96 mA = 8.96*10^-3 A